# A Survey of Artificial Intelligence in Gait-Based Neurodegenerative Disease Diagnosis

### v2

@@ 4.2.2 C-DL Models for PD Diagnosis
[125] (ID = P-37) proposes a Locally Weighted Random Forest (LWRF) regression model to predict the PD symptoms using time- and frequency-domain features derived from vGRF signals collected using foot worn sensors. [126] (ID = P-39) develop a hybrid model combining Long Short-Term Memory (LSTM) and Convolutional Neural Network (CNN) for PD’s detection and severity rating. The model captures the spatio-temporal characteristics of gait dynamics using vGRF signals and offers a multi-category classification approach for PD severity, beyond the binary detection typically provided by other ML methods. The study of [81] (ID = P-43) leverages convolutional LSTM to identify PD patients from Kinect-based skeleton data and achieves up to 83% accuracy. It also suggests that the data preprocessing ($e.g.$, cropped data) can help boost the classification performance. [127] (ID = P-52) devises a deep neural network (DNN) based model to identify Parkinsonian gait based on vGRF data and achieves 98.7% accuracy. This is also the first study to predict the Unified Parkinson’s Disease Rating Scale (UPDRS) severity ($i.e.$, Parkinsion severity rate) of a subject and achieve 85.3% accuracy. [128] (ID = P-58) utilizes both CNN and attention-enhanced LSTM networks for left and right gait analysis, which helps achieve high-competitive PD classification accuracy (98.03%). [129] (ID = P-63) uses balance and gait variables obtained with wearable inertial motion sensors to differentiate PD patients using machine learning models (NN, SVM, KNN, DT, RF). [130] (ID = P-64) applies a mask R-CNN model for extracting human silhouettes from video frames, and then obtains gait energy images (GEIs) from human silhouettes as features for PD gait classification. [131] (ID = P-66) presents a novel Bidirectional Long Short-Term Memory (Bi-LSTM) model to classify PD from HC using motion data, acquired during performing 13 tasks derived from the Movement Disorder Society-Unified Parkinson’s Disease Rating Scale III (MDS-UPDRS III), with wearable sensors. [132] (ID = P-67) reports a smart-phone based PD diagnosis method by utilizing real-world accelerometer records and achieved an AUC score of 0.86. This method demonstrate the superior accuracy for at-home screening of PD and other NDD conditions affecting movement while acknowledges limitations such as potential over-enrollment of tremor-dominant PD subjects in their experiment. [133] (ID = P-68) uses a novel technique called Neighborhood Representation Local Binary Pattern (NR-LBP) and ANN to model the relationship between statistical features extracted from vGRF data and PD symptoms. [134] (ID = P-69) utilizes LSTM to differentiate the PD patients and healthy subjects based on the vGRF and gait time series, which performs better than other state-of-the-art methods with binary-classification accuracy of 98.6%. The Bi-LSTM model is leveraged in [135] (ID = P-70) to classify PD gait using vGRF signals from single sensors. This study reveals that combining time-frequency and time-space features of physiological signals are more effective and can boost the classification accuracy of deep-learning models. In [136] (ID = P-74), force domain, peak domain, and abnormality domain features (22 features) are combined to identify PD patients based on CNN and XGBoost, with an accuracy of 95.68%-97.32%. [84] (ID = P-76) characterizes the pre-diagnostic period of PD and develops a machine learning model (LR) and deep learning model (RNN) for PD diagnosis. In [137] (ID = P-84), the authors focus on identifying changes in gait patterns associated with PD by analyzing the spatio-temporal data of vGRF. They propose a model based on CNN that evaluates the changes in vGRF and identify the Parkinsonian gait from healthy gait with an accuracy of 97.0%. [63] (ID = P-86) combines CNN and gated recurrent units (GRUs) to perform classification between PD and health subjects from IMU signals, achieving 83.7-92.7% accuracy. The results suggest that the combination of temporal and spectral information is more effective in classifying gait patterns of PD patients. [138] (ID = P-104) devises an anomaly detection algorithm to classify PD and healthy controls based on wrist-sensor recordings (tri-axial accelerometer data) and achieves 83.3% accuracy. [139] (ID = P-115) suggests a real-time PD identification method using wavelet transform-based features extracted from vGRF data. Experimental results using NN with resilient back-propagation algorithm achieved an average efficiency of 86.75%. 2D CNN is leveraged in [140] (ID = P-116) to diagnose PD at an early stage from vGRF signal data and achieves 98% accuracy. In [141] (ID = P-120), a plantar pressure analysis method is proposed to identify the gait dynamics of PD patients using a constant radial basis function neural network (RBFNN).
@@ 4.2.3 A-DL Models for PD Diagnosis
In [142] (ID = P-15), a novel hybrid model based on BP neural networks is proposed to classify PD and HC, which outperforms four baseline methods on a public dataset with an average accuracy of 87.9%. [143] (ID = P-17) introduces a Q-backpropagated Time Delay Neural Network (Q-BTDNN), which harvests the temporal gait data collected from wearable devices, combined with the Q-backpropagation (Q-BP) algorithm for predicting gait disturbances in PD. The study of [144] (ID = P-31) proposes a computational method based on tensor decomposition for analyzing gait dynamics in PD from multisensor time series data, achieving an accuracy of 86.2%. [145] (ID = P-75) uses deep CNN to classify gait deterioration in PD patients based on spatio-temporal vGRF signals and achieves 95.5% accuracy in average. The study suggests that the heel strike and body balance are the most indicative gait elements in PD classification. [146] (ID = P-82) develops an NN model to classify PD and HC based on the vGRF dataset collected from a public dataset and achieves an accuracy result of 84.8%. [147] (ID = P-85) proposes an interval type-2 fuzzy neural network with novel quasi-Levenberg-Marquardt (qLM) learning approach to identify PD patients from the gait cycle based vGRF signals. Based on the 10 clinical features extracted from 14 gait features, the model achieves 97.6% accuracy with a smaller size of parameters and a good interpretability ($i.e.$, rules based on fuzzy logic). In [148] (ID = P-93), the authors propose a novel architecture PD-ResNet based on ResNet [95] and use it to differentiate PD and HC. Its result outperforms many existing state-of-the-art models with 95.5% accuracy. A hybrid deep learning model combining CNN and locally weighted RF (LWRF) [149] (ID = P-97) is devised to perform PD detection tasks from time and frequency features of vGRF, achieving 99.5% accuracy. [150] (ID = P-98) presents a Riemannian deep learning strategy called SPDnet to learn compact descriptors for discriminating Parkinsonian gait patterns. The proposed SPDnet learns discriminatory patterns between Parkinson’s and control populations, outperforming traditional 3D convolutional representations with 99.4% accuracy. [77] (ID = P-106) proposes a correlative memory neural network (CorrMNN) scheme to extract spatio-temporal features to discriminate NDDs’ gait from normal gait based on vGRF, and achieves up to 100% accuracy. A video-based SlowFast GCN network is proposed in [151] (ID = P-109) to leverage 2D pose body-joint graph for diagnosing gait disorders such as PD. [152] (ID = P-121) proposes a novel model named WM-STGCN that leverages a Weighted adjacency matrix and Multi-scale temporal convolution in a spatio-temporal Graph Convolution Network to classify PD gaits from normal gaits and achieves up to 87.1% accuracy on 2D skeleton data. This method outperforms many existing methods including LSTM, KNN, DT, AdaBoost, and ST-GCN models [153]. [154] (ID = P-123) devises a static-dynamic temporal networks that consists of one-dimensional, two-dimensional convolutional networks, and attention mechanisms to classify PD patients based on foot force (vGRF time series) and achieves 96.7% accuracy.
@@
Table 2: Existing studies (P-1 to P-124) of AI-assisted PD diagnosis based on gait. We summarize input data, main AI model(s), experimental sample size, and accuracy (range) of each study. Quality scores ($S_{Q}$) that evaluate advancement of AI models, comprehensiveness of comparison, and sufficiency of samples are computed (see Sec. 4.1).
@@
P-47 [116] 2019 PD Sensor Force sensitive resistor SVM 166 88.9-100.0 0.78
@@
P-53 [66] 2020 PD Sensor IMUs+Force sensitive resistor SVM 10 97.4-98.8 0.33
P-54 [176] 2020 PD Sensor vGRF Discriminant analysis 47 95.0 0.44
@@
P-61 [179] 2020 PD Sensor IMUs PLS-DA 93 97.9 0.44
P-62 [180] 2020 PD Sensor IMUs LR 16 93.6 0.33
@@
Table 4: Continued table of 3 for AI-assisted PD diagnosis based on gait.
@@
P-79 [183] 2021 PD Sensor vGRF LSTM 11 84.5 0.67
P-80 [184] 2021 PD Vision Camera CNN 22 94.9 0.44
@@
P-92 [82] 2022 PD Vision Camera (Kinect) LR, SVM, DT, NB, RF 60 81.8-84.5 0.44
@@
P-94 [190] 2022 PD Sensor Force sensitive resistor SVM, RF 22 99.6 0.33
@@
22 99.4 0.67
P-99 [193] 2022 PD Sensor IMUs CNN, NN 17 91.0-92.0 0.56
Table 5: Continued table of 4 for AI-assisted PD diagnosis based on gait.
@@
P-102 [78] 2022 PD Sensor+Vision Camera+Force sensitive resistor AdaBoost, Extra Trees, RF 44 88.0 0.44
@@
P-110 [68] 2022 PD Sensor Force sensitive resistor+IMUs RF, LR, Gradient Boosting 22 97.3 0.33
@@
30 94.0-96.0 0.44
P-118 [202] 2023 PD Sensor IMUs CNN 90 82.8-86.4 0.55
P-119 [203] 2023 PD Sensor IMUs PCA, SVM 280 88.0-90.0 0.55
@@
P-124 [205] 2023 PD Sensor IMUs Non-linear model 53 84.9 0.44
Table 6: Existing studies (C-1 to C-19) that include combinatorial diagnosis of different NDs (PD, ALS, HD) based on gait. We summarize input data, main AI model(s), experimental sample size, accuracy (range), and quality score of each study. Quality scores that evaluate the novelty of AI models, comprehensiveness of method comparison, and sufficiency of experimental samples are computed (see Sec. 4.1).
@@
C-1 [101] 2012 PD, ALS, HD Sensor Force sensitive resistor SVM 64 90.6 0.44
@@
C-3 [102] 2015 PD, ALS, HD Sensor Force sensitive resistor SVM, RF, MLP, KNN 64 96.6-100.0 0.44
C-4 [207] 2017 PD, ALS, HD Sensor vGRF SVM, EC, DT, KNN 62 79.3 0.56
C-5 [208] 2017 PD, ALS, HD Sensor vGRF Adaboost 64 87.3-92.3 0.67
C-6 [209] 2018 PD, ALS, HD Sensor vGRF RF, SVM 64 90.6-96.9 0.44
C-7 [210] 2018 PD, ALS, HD Sensor Force sensitive resistor fusszy KNN 64 98.4 0.44
C-8 [211] 2020 PD, ALS, HD Sensor Force sensitive resistor Random forest 64 79.3-91.7 0.67
C-9 [212] 2020 PD, ALS, HD Vision Camera Discriminant Analysis, RF, Multi-SVM, KNN 64 92.2 0.44
C-10 [213] 2020 PD, ALS, HD Sensor Force sensitive resistor SVM, KNN 64 99.8-100.0 0.44
C-11 [214] 2020 PD, ALS, HD Sensor vGRF CNN 64 98.4-100.0 0.56
@@
64 99.8-100.0 0.89
C-13 [215] 2020 PD, ALS, HD Sensor Force sensitive resistor KNN 64 90.5 0.44
C-14 [216] 2021 PD, ALS, HD Sensor vGRF CNN, SVM 64 97.4-100.0 0.78
@@
C-17 [218] 2022 PD, ALS, HD Sensor vGRF RF, classification trees 64 69.6-95.1 0.67
@@
Table 7: Existing studies of AI-assisted diagnosis of AD (A-1 to A-5), ALS (S-1 to S-3), HD (H-1 to H-7) or MS (M-1 to M-6) based on gait. We summarize input data, main AI model(s), experimental sample size, accuracy (range), and quality score of each study. Quality scores that evaluate the novelty of AI models, comprehensiveness of method comparison, and sufficiency of experimental samples are computed (see Sec. 4.1).
@@
A-1 [221] 2016 AD Sensor IMUs SVM, MLP, RBNs, DBSs 72 91.0-96.6 0.56
@@
A-3 [222] 2020 AD Sensor Pressure sentsitive mat SVM 78 78.0 0.44
A-4 [107] 2022 AD Sensor+Vision Kinect depth sensor + Camera SVM 85 97.8 0.44
@@
S-1 [104] 2016 ALS Sensor Force sensitive resistor SVM 28 92.9 0.33
S-2 [224] 2020 ALS Sensor Force sensitive resistor SVM, KNN. DT 29 86.2-96.6 0.33
S-3 [225] 2021 ALS Sensor Force sensitive resistor SVM, KNN, NB, LDA, etc 29 89.7 0.44
H-1 [48] 2016 HD Sensor IMUs HMM, SVM 42 88.2 0.44
H-2 [226] 2018 HD Sensor IMUs DT, RF, Logitboost, Multiboost 14 78.6-92.9 0.56
H-3 [227] 2019 HD Sensor Walkway system SVM, 3D CNN 12 82.0-86.9 0.56
H-4 [228] 2019 HD Sensor Force sensitive resistor SVM, KNN, NB, LDA, etc 36 58.3-94.4 0.44
H-5 [229] 2019 HD Sensor Force sensitive resistor SVM, KNN, NB, LDA, etc 36 100.0 0.67
H-6 [230] 2019 HD Sensor vGRF SVM, NB, DT, RF, etc 35 85.7-97.1 0.44
@@
M-2 [59] 2021 MS Sensor Treadmill SVM, RF, AdaBoost, MLP, etc 40 94.3 0.44
M-3 [232] 2021 MS Sensor Force sensitive sensor SVM, Gaussian Naive Bayes, DT, KNN, etc 60 74.5 0.44
@@
M-5 [106] 2021 MS Sensor IMUs SVM, LR, RF 97 79.3-85.1 0.44
M-6 [58] 2022 MS Sensor Walkway system SVM, LR, XGBoost 88 77.0-81.0 0.44

### v3

@@ 4.2.2 C-DL Models for PD Diagnosis
Acsurouglu et al. [138] (ID = P-37) proposed a Locally Weighted Random Forest (LWRF) regression model to predict the PD symptoms using time- and frequency-domain features derived from vGRF signals collected using foot worn sensors. Zhao et al. [139] (ID = P-39) developed a hybrid model combining Long Short-Term Memory and Convolutional Neural Network for PD’s detection and severity rating. The model captured the spatio-temporal characteristics of gait dynamics using vGRF signals and offered a multi-category classification approach for PD severity, beyond the binary detection typically provided by other ML methods. The study by Reyes et al. [86] (ID = P-43) leveraged convolutional LSTM to identify PD patients from Kinect-based skeleton data and achieved up to 83% accuracy. It also suggested that the data preprocessing ($e.g.$, cropped data) could help boost the classification performance. El et al. [140] (ID = P-52) devised a deep neural network (DNN) based model to identify Parkinsonian gait based on vGRF data and achieved 98.7% accuracy. This was also the first study to predict the Unified Parkinson’s Disease Rating Scale (UPDRS) severity ($i.e.$, Parkinsion severity rate) of a subject and achieved a 85.3% accuracy. Xia et al. [141] (ID = P-58) utilized both CNN and attention-enhanced LSTM networks for left and right gait analysis, which helps achieved highly competitive PD classification accuracy (98.03%). Moon et al. [142] (ID = P-63) used balance and gait variables obtained with wearable inertial motion sensors to differentiate PD patients using machine learning models (NN, SVM, KNN, DT, RF). Gong et al. [143] (ID = P-64) applied a mask R-CNN model for extracting human silhouettes from video frames, and then obtained gait energy images (GEIs) from human silhouettes as features for PD gait classification. Butt et al. [144] (ID = P-66) presented a novel Bidirectional Long Short-Term Memory (Bi-LSTM) model to classify PD from HC using motion data, acquired during performing 13 tasks derived from the Movement Disorder Society-Unified Parkinson’s Disease Rating Scale III (MDS-UPDRS III), with wearable sensors. Zhang et al. [145] (ID = P-67) reported a smart-phone based PD diagnosis method by utilizing real-world accelerometer records and achieved an AUC score of 0.86. This method demonstrated the superior accuracy for at-home screening of PD and other NDD (Neurodegenerative Disease) conditions affecting movement while acknowledges limitations such as potential over-enrollment of tremor-dominant PD subjects in their experiment. Yurdakul et al. [146] (ID = P-68) used a novel technique called Neighborhood Representation Local Binary Pattern (NR-LBP) and ANN to model the relationship between statistical features extracted from vGRF data and PD symptoms. Balaji et al. [147] (ID = P-69) utilized LSTM to differentiate the PD patients and healthy subjects based on the vGRF and gait time series, which performed better than other state-of-the-art methods with binary-classification accuracy of 98.6%. The Bi-LSTM model was leveraged in the work by Pham et al. [148] (ID = P-70) to classify PD gait using vGRF signals from single sensors. This study revealed that combining time-frequency and time-space features of physiological signals were more effective and could boost the classification accuracy of deep-learning models. Force domain, peak domain, and abnormality domain features (22 features) were combined to identify PD patients based on CNN and XGBoost, with an accuracy of 95.68%-97.32% (Ma et al. [149] (ID = P-74)). Yuan et al. [89] (ID = P-76) characterized the pre-diagnostic period of PD and develops a ML model (i.e., LR) and DL model (i.e., RNN) for PD diagnosis. The authors focused on identifying changes in gait patterns associated with PD by analyzing the spatio-temporal data of vGRF. They proposed a model based on CNN that evaluated the changes in vGRF and identify the Parkinsonian gait from healthy gait with an accuracy of 97.0% (Ahamed et al. [150] (ID = P-84)). Carvajal et al. [68] (ID = P-86) combined CNN and gated recurrent units (GRUs) to perform classification between PD and health subjects from IMU signals, achieving a 83.7-92.7% accuracy. The results suggested that the combination of temporal and spectral information was more effective in classifying gait patterns of PD patients. Brand et al. [151] (ID = P-104) devised an anomaly detection algorithm to classify PD and healthy controls based on wrist-sensor recordings (tri-axial accelerometer data) and achieved a 83.3% accuracy. Baby et al. [152] (ID = P-115) suggested a real-time PD identification method using wavelet transform-based features extracted from vGRF data. Experimental results using NN with resilient back-propagation algorithm achieved an average efficiency of 86.75%. 2D CNN was leveraged by Divyashree et al. [153] (ID = P-116) to diagnose PD at an early stage from vGRF signal data and achieved 98% accuracy. Sun et al. [154] (ID = P-120) proposed a plantar pressure analysis method to identify the gait dynamics of PD patients using a constant radial basis function neural network (RBFNN).
@@ 4.2.3 A-DL Models for PD Diagnosis
In the work of Wang et al.[108] (ID = P-15), a novel hybrid model based on BP neural networks was proposed to classify PD and HC, which outperformed four baseline methods on a public dataset with an average accuracy of 87.9%. Jane [155] (ID = P-17) introduced a Q-backpropagated Time Delay Neural Network (Q-BTDNN), which harvested the temporal gait data collected from wearable devices, combined with the Q-backpropagation (Q-BP) algorithm for predicting gait disturbances in PD. The study of Pham et al. [156] (ID = P-31) proposed a computational method based on tensor decomposition for analyzing gait dynamics in PD from multi-sensor time series data, achieving an accuracy of 86.2%. Alharthi et al. [157] (ID = P-75) used deep CNN to classify gait deterioration in PD patients based on spatio-temporal vGRF signals and achieves 95.5% accuracy in average. The study suggested that the heel strike and body balance are the most indicative gait elements in PD classification. Nguyen et al. [158] (ID = P-82) developed an NN model to classify PD and HC based on the vGRF dataset collected from a public dataset and achieved an accuracy result of 84.8%. Salmi et al. [159] (ID = P-85) proposed an interval type-2 fuzzy neural network with novel quasi-Levenberg-Marquardt (qLM) learning approach to identify PD patients from the gait cycle based vGRF signals. Based on the 10 clinical features extracted from 14 gait features, the model achieved a 97.6% accuracy with a smaller size of parameters and a good interpretability ($i.e.$, rules based on fuzzy logic). In the work by Yang et al. [160] (ID = P-93), the authors proposed a novel architecture PD-ResNet based on ResNet [100] and used it to differentiate PD and HC. Their result outperformed many existing state-of-the-art models with 95.5% accuracy. A hybrid deep learning model combining CNN and locally weighted RF (LWRF) was devised to perform PD detection tasks from time and frequency features of vGRF, achieving a 99.5% accuracy (Acsurouglu et al. [161] (ID = P-97). Zhao et al. [82] (ID = P-106) proposed a correlative memory neural network (CorrMNN) scheme to extract spatio-temporal features to discriminate NDs’ gait from normal gait based on vGRF, and achieved up to 100% accuracy. A video-based SlowFast GCN network was proposed by Zeng et al. [162] (ID = P-109) to leverage 2D pose body-joint graph for diagnosing gait disorders such as PD. Zhang et al. [111] (ID = P-121) proposed a novel model named WM-STGCN that leveraged a weighted adjacency matrix and multi-scale temporal convolution in a spatio-temporal graph convolution network to classify PD gaits from normal gaits and achieved up to 87.1% accuracy on 2D skeleton data. This method outperformed many existing methods including LSTM, KNN, DT, AdaBoost, and spatial temporal graph convolutional networks (ST-GCN) model by Yan et al. [163]. Dong et al.[109] (ID = P-123) devised a static-dynamic temporal networks that consisted of one-dimensional, two-dimensional convolutional networks, and attention mechanisms to classify PD patients based on foot force (vGRF time series) and achieved a 96.7% accuracy. The necessity for objective and efficient gait assessment in PD using the Movement Disorder Society Unified Parkinson’s Disease Rating Scale (MDS-UPDRS) is proposed in [164] (ID = P-126, $S_{Q}=1$).To overcome the limitations of existing PD gait skeleton datasets, they collect a sizable dataset including 102 PD patients, 30 healthy older adults, and 16 young adults using multi-view Azure Kinect sensors. In addition, this paper develops a cross-spatiotemporal graph convolution network (CST-GCN) that captures complex gait features by modeling cross-spacetime dependencies between joints to assess age-related effects on PD gait.
@@
Table 2: Existing studies (P-1 to P-126) of AI-assisted PD diagnosis based on gait. We summarize input data, main AI model(s), experimental sample size, and accuracy (range) of each study. Quality scores ($S_{Q}$) that evaluate advancement of AI models, comprehensiveness of comparison, and sufficiency of samples are computed (see Sec. 4.1).
@@
P-47 [andrei2019parkinso] 2019 PD Sensor Force sensitive resistor SVM 166 88.9-100.0 0.78
@@
P-53 [71] 2020 PD Sensor IMUs+Force sensitive resistor SVM 10 97.4-98.8 0.44
P-54 [186] 2020 PD Sensor vGRF Discriminant analysis 47 95.0 0.56
@@
P-61 [189] 2020 PD Sensor IMUs PLS-DA 93 97.9 0.56
P-62 [190] 2020 PD Sensor IMUs LR 16 93.6 0.44
@@
P-79 [193] 2021 PD Sensor vGRF LSTM 11 84.5 0.56
P-80 [194] 2021 PD Vision Camera CNN 22 94.9 0.33
@@
P-92 [87] 2022 PD Vision Camera (Kinect) LR, SVM, DT, NB, RF 60 81.8-84.5 0.56
@@
P-94 [200] 2022 PD Sensor Force sensitive resistor SVM, RF 22 99.6 0.44
@@
22 99.4 0.56
@@
P-102 [83] 2022 PD Sensor+Vision Camera+Force sensitive resistor AdaBoost, Extra Trees, RF 44 88.0 0.56
@@
P-110 [73] 2022 PD Sensor Force sensitive resistor+IMUs RF, LR, Gradient Boosting 22 97.3 0.44
@@
30 94.0-96.0 0.56
P-118 [212] 2023 PD Sensor IMUs CNN 90 82.8-86.4 0.56
P-119 [213] 2023 PD Sensor IMUs PCA, SVM 280 88.0-90.0 0.56
@@
P-124 [214] 2023 PD Sensor IMUs Non-linear model 53 84.9 0.56
P-125 [137] 2024 PD Sensor Wearable devices Random forest 132 92.3 0.88
P-126 [164] 2023 PD Sensor Kinect
Cross-spatiotemporal graph
convolution network (CST-GCN)
148 84.9 1.00
Table 6: Existing studies (C-1 to C-21) that include combinatorial diagnosis of different NDs (PD, ALS, HD) based on gait. We summarize input data, main AI model(s), experimental sample size, accuracy (range), and quality score of each study. Quality scores that evaluate the novelty of AI models, comprehensiveness of method comparison, and sufficiency of experimental samples are computed (see Sec. 4.1).
@@
C-1 [112] 2012 PD, ALS, HD Sensor Force sensitive resistor SVM 64 90.6 0.56
@@
C-3 [113] 2015 PD, ALS, HD Sensor Force sensitive resistor SVM, RF, MLP, KNN 64 96.6-100.0 0.56
C-4 [216] 2017 PD, ALS, HD Sensor vGRF SVM, EC, DT, KNN 62 79.3 0.67
C-5 [217] 2017 PD, ALS, HD Sensor vGRF Adaboost 64 87.3-92.3 0.78
C-6 [218] 2018 PD, ALS, HD Sensor vGRF RF, SVM 64 90.6-96.9 0.56
C-7 [219] 2018 PD, ALS, HD Sensor Force sensitive resistor fusszy KNN 64 98.4 0.56
C-8 [220] 2020 PD, ALS, HD Sensor Force sensitive resistor Random forest 64 79.3-91.7 0.78
C-9 [221] 2020 PD, ALS, HD Vision Camera Discriminant Analysis, RF, Multi-SVM, KNN 64 92.2 0.56
C-10 [222] 2020 PD, ALS, HD Sensor Force sensitive resistor SVM, KNN 64 99.8-100.0 0.56
C-11 [223] 2020 PD, ALS, HD Sensor vGRF CNN 64 98.4-100.0 0.67
@@
64 99.6-100.0 0.89
C-13 [224] 2020 PD, ALS, HD Sensor Force sensitive resistor KNN 64 90.5 0.56
C-14 [225] 2021 PD, ALS, HD Sensor vGRF CNN, SVM 64 97.4-100 0.78
@@
C-17 [227] 2022 PD, ALS, HD Sensor vGRF RF, classification trees 64 69.6-95.1 0.78
@@
C-20 [230] 2024 PD, ALS, HD Sensor Force sensitive resistor CNN 64 93.56-97.65 0.77
C-21 [231] 2024 PD, ALS, HD Sensor Force sensitive resistor CNN-BiLSTM 64 95.43-1 1.00
Table 7: Existing studies of AI-assisted diagnosis of AD (A-1 to A-6), ALS (S-1 to S-3), HD (H-1 to H-7) or MS (M-1 to M-6) based on gait. We summarize input data, main AI model(s), experimental sample size, accuracy (range), and quality score of each study. Quality scores that evaluate the novelty of AI models, comprehensiveness of method comparison, and sufficiency of experimental samples are computed (see Sec. 4.1).
@@
A-1 [232] 2016 AD Sensor IMUs SVM, MLP, RBNs, DBSs 72 91.0-96.6 0.67
@@
A-3 [233] 2020 AD Sensor Pressure sentsitive mat SVM 78 78.0 0.56
A-4 [118] 2022 AD Sensor+Vision Kinect depth sensor + Camera SVM 85 97.8 0.56
@@
A-6 [235] 2024 AD Sensor Accelerometer Sensor (Smartphone) CLADSI 35 84-87 0.77
S-1 [115] 2016 ALS Sensor Force sensitive resistor SVM 28 92.9 0.44
S-2 [236] 2020 ALS Sensor Force sensitive resistor SVM, KNN. DT 29 86.2-96.6 0.44
S-3 [237] 2021 ALS Sensor Force sensitive resistor SVM, KNN, NB, LDA, etc 29 89.7 0.67
H-1 [53] 2016 HD Sensor IMUs HMM, SVM 42 88.2 0.56
H-2 [238] 2018 HD Sensor IMUs DT, RF, Logitboost, Multiboost 14 78.6-92.9 0.67
H-3 [239] 2019 HD Sensor Walkway system SVM, 3D CNN 12 82.0-86.9 0.44
H-4 [240] 2019 HD Sensor force sensitive resistor SVM, KNN, NB, LDA, DT 36 58.3-94.4 0.56
H-5 [241] 2019 HD Sensor force sensitive resistor SVM, KNN, NB, LDA, DT 36 100.0 0.78
H-6 [242] 2019 HD Sensor vGRF SVM, NB, DT, RF, LR 35 85.7-97.1 0.56
@@
M-2 [64] 2021 MS Sensor Treadmill SVM, RF, AdaBoost, MLP, etc 40 94.3 0.56
M-3 [244] 2021 MS Sensor Force sensitive sensor SVM, Gaussian Naive Bayes, DT, KNN, etc 60 74.5 0.56
@@
M-5 [117] 2021 MS Sensor IMUs SVM, LR, RF 97 79.3-85.1 0.56
M-6 [63] 2022 MS Sensor Walkway system SVM, LR, XGBoost 88 77.0-81.0 0.56